# Refereed Learning

### v1

@@ Abstract
We initiate an investigation of learning tasks in a setting where the learner is given access to two competing provers, only one of which is honest. Specifically, we consider the power of such learners in assessing purported properties of opaque models. Following prior work that considers the power of competing provers in different settings, we call this setting refereed learning.
After formulating a general definition of refereed learning tasks, we show refereed learning protocols that obtain a level of accuracy that far exceeds what is obtainable at comparable cost without provers, or even with a single prover. We concentrate on the task of choosing the better one out of two black-box models, with respect to some ground truth. While we consider a range of parameters, perhaps our most notable result is in the high-precision range: For all $\varepsilon>0$ and ambient dimension $d$, our learner makes only one query to the ground truth function, communicates only $(1+\frac{1}{\varepsilon^{2}})\cdot\text{poly}(d)$ bits with the provers, and outputs a model whose loss is within a multiplicative factor of $(1+\varepsilon)$ of the best model’s loss. Obtaining comparable loss with a single prover would require the learner to access the ground truth at almost all of the points in the domain. To obtain this bound, we develop a technique that allows the learner to sample, using the provers, from a distribution that is not efficiently samplable to begin with. We find this technique to be of independent interest.
@@ 1 Introduction
Modern machine learning tasks require increasingly large amounts of data and computational power. As a result, model training has shifted from a task that almost anyone can perform on their own to a task that requires the assistance of external agents that are resource-abundant and have better access to the underlying data. Furthermore, one is often presented with opaque models that purport to approximate some ground truth, but are not accompanied with a rigorous or trustworthy performance guarantee. Moreover, such models are often given only as black-boxes via a controlled query/response mechanism, and without disclosing their oft-extensive training processes.
@@
This state of affairs naturally raises the need to verify claims of performance of such models, without fully trusting the parties making the claims, and with significantly fewer resources than those needed to train comparable models to begin with. Verifying such claims appears hard: There is a rich literature on efficient verification of computations performed by powerful-but-untrusted parties (e.g., [Kil92, Mic94, GKR15]; see [WB15] for a survey); however, such mechanisms appear to be ill-suited for the task of verifying the performance of even explicitly described ML models, let alone opaque ones. There is also a growing body of work on using powerful-but-untrusted intermediaries to learn properties of unknown ground-truth functions (e.g., [GRSY21, CK21]), as well as efficient verification of claims made by powerful provers regarding properties of huge combinatorial objects in general [EKR04, RVW13]. However, these works mostly focus on verifiable execution of a specific learning (or property testing) algorithm, rather than evaluating the performance of a given, opaque model without knowledge of the process used to train it. Perhaps closest to our task are the works of Herman and Rothblum [HR22, HR24b, HR24a] that allow verifying claims about properties of distributions that are accessible only via obtaining samples. However, this is still a far cry from assessing properties of black-box models.
@@
Some natural properties of black-box models can of course be assessed using standard methods. For instance, the loss of a given model w.r.t. some sample distribution and loss function can be approximated by computing the empirical loss on a large enough sample. However, this method can be prohibitively costly both in samples from the ground truth and in queries to the model. One can use a technique from [GRSY21] to push much of the burden to an external powerful-but-untrusted prover, but even this method incurs high cost: to obtain an additive bound of $\eta$ on the error, the learner needs to both communicate $\eta^{-2}$ unlabeled samples to the prover and have the prover query and report the ground truth values at these points, and obtain $\eta^{-1}$ labeled samples which are hidden from the prover in order to verify the prover’s responses.
@@
We would like to do better—in terms of the error, in terms of the access to the ground truth, and in terms of the communication with the prover. However, as evidenced by lower bounds proven in [GRSY21], this appears hard—at least within the present framing of the problem.
@@ 1.1 This work
We show that the quality of black-box models can be assessed with significantly better accuracy and with significantly lower cost, if the learner can interact with two powerful and competing provers, one of which is honest. The provers’ power can be manifested either in terms of their computational power, or in access to the ground truth, or in knowledge of the models, or any combination of these. This model can be naturally viewed as an extension of the refereed delegation of computation model [FST88, FK97, CRR11, CRR13, KR14] to our setting. Following the cue of these works, we coin the term refereed learning to denote the type of learning performed in this model. (See Section 1.2 for more discussion on refereed delegation and other related works.)
We first define refereed learning for a general setting. Next we focus on applying refereed learning to the following specific task, where we showcase the power of the refereed learning framework via concrete protocols. The learner-verifier [^1] We use the terms learner and verifier interchangeably. Indeed, the learner now doubles as a verifier. is presented with two candidate models that purport to compute the same ground truth, and is tasked with choosing the model that incurs the smaller overall loss with respect to some sample distribution and loss function. (The restriction to two candidates is not essential, but it helps make the model more concrete. It also facilitates envisioning each one of the provers as “trying to promote” a different one of the two models.)
The salient parameters we consider are (a) the overall loss of the output model relative to the better of the two competing ones; (b) the number of learner queries to the ground truth and samples from it (both labeled and unlabeled ones); (c) the number of learner queries to the candidate models; (d) the computational complexity of the learner; (e) the computational complexity and query and sample complexity of the provers. We first sketch and briefly discuss definitions of refereed learning, then present our results, and finally discuss the new tools we develop to obtain these results.
@@ 1.1.1 Defining refereed learning
We start with general refereed learning. Consider a learner-verifier $\mathcal{V}$ and two provers ${\mathcal{P}}_{0},{\mathcal{P}}_{1}$ that are presented with a ground truth function $f:\{0,1\}^{d}\rightarrow\mathcal{Y}$, a distribution $\mathcal{D}$} over $\{0,1\}^{d}$, and $k$ models (a.k.a. hypotheses) $h_{1},\ldots,h_{k}\in{\mathcal{H}}$, where $\mathcal{H}$} is some family of hypotheses. We will typically assume that $h_{1},\ldots,h_{k}$ and $f$ are accessed via queries and, depending on the setting, that $\mathcal{D}$} is accessed either via samples or via queries to its probability mass function, $Q_{\mathcal{D}}$, which maps $x\mapsto\operatorname*{{Pr}}_{X\sim{\mathcal{D}}}{\left[{X=x}\right]}$. To measure the learner’s performance we use a scoring function $\mathcal{S}$ (which assigns a score to each potential output of the learner, with respect to some sample distribution $\mathcal{D}$}, hypotheses $h_{1},\ldots,h_{k}$ and function $f$) and a target function $\mathcal{T}$}.
@@
A protocol ${\mathcal{P}}_{0},{\mathcal{P}}_{1},\mathcal{V}]$ is an $(\alpha,\eta,\beta)$-refereed learning protocol, with respect to a family $\mathcal{H}$} of hypotheses, a scoring function $\mathcal{S}$ and target function $\mathcal{T}$}, if for all $b\in\{0,1\}$, $h_{1},\ldots,h_{k}\in{\mathcal{H}}$, and ${\mathcal{P}}^{*}_{1-b}$, the learner output $\rho$ satisfies
@@
In this general definition, we choose to minimize the score $\mathcal{S}$. This choice is motivated by the learning theory formulation of this problem, used in the specific definition below, where the goal is to select the model (hypothesis) that minimizes a loss function quantifying the deviation of the hypothesis from some ground truth. Concretely, if the learner is restricted to selecting between two models (hypotheses) $h_{0}$ and $h_{1}$, its output bit $\rho$ indicates the selected hypothesis, the general scoring function $\mathcal{S}$ measures the loss of the chosen model with respect to some metric $\ell$ on the domain $\mathcal{Y}$, and the target function $\mathcal{T}$} is the smaller of the losses of $h_{0}$ and $h_{1}$. (Specifically, we let the loss of some hypothesis $h$ w.r.t. sample distribution $\mathcal{D}$} and ground truth $f$ be $\mathcal{L}_{{\mathcal{D}}}{\left({f,h\mid\ell}\right)}=\mathop{{\mathds{E}}\/}_{x\sim{\mathcal{D}}}{\left[{\ell{\big({f(x),h(x)}\big)}}\right]}$.) That is:
@@
An alternative formulation would instead bound the expected score (resp. loss). That is, the requirement would be that $\mathop{{\mathds{E}}\/}[\mathcal{S}(\rho,{\mathcal{D}},f,h_{1},\ldots,h_{k}))]\leq\alpha{\mathcal{T}}({\mathcal{D}},f,h_{1},\ldots,h_{k})+\eta$ (in the general case), or $\mathop{{\mathds{E}}\/}[\mathcal{L}_{\mathcal{D}}(f,h_{\rho}\mid\ell)]\leq\alpha\min_{s\in\{0,1\}}\mathcal{L}_{\mathcal{D}}(f,h_{s}\mid\ell)+\eta$ (in the case of minimizing the loss). While the two formulations are incomparable in general, our protocols satisfy both with similar parameters.
@@
The above definition posits that at least one of the provers follows the protocol. We note, however, that refereed learning protocols appear to preserve their guarantees even when both provers are strategic with opposing goals. Some supporting evidence for the implication is the use of protocols developed in the refereed delegation of computation model in real-world applications where truth-telling is economically incentivized. Similar phenomena are manifested in the context of debate systems. See more details in Section 1.2.
@@ 1.1.2 Our results
We give protocols for both the additive and multiplicative error settings. In the additive error setting, we show how to use the two provers to obtain additive error similar to that obtained by the [GRSY21] protocol mentioned above, while significantly reducing the learner’s interaction with both the ground truth and the models: this interaction now consists of only a single query.
@@
Next we design protocols for the low-loss setting, which turns out to be significantly more challenging. Here we would like to guarantee that the learner makes the right choice even when the models’ losses are close to each other, up to a multiplicative factor of $1+\varepsilon$ for an arbitrarily small $\varepsilon>0$. Indeed, in such a setting, the number of samples needed to even observe the difference can be close to the entire sample space. Still, determining which of the two competing models incurs smaller loss may have significant ramifications in applications that require high precision (e.g., using ML models for medical predictions based on imaging or other multi-dimensional measurements, or for financial applications where even tiny error margins become significant over time).
@@
We then extend these results to handle arbitrary metric loss functions as well as arbitrary sample distributions. Our results also take into account the cost of obtaining the desired level of numerical precision. The protocol for this more general setting is guaranteed to select a model whose loss is at most a multiplicative factor of $3+\varepsilon$ worse than the best model. Moreover, the learner makes a single query to either $f$, the two models, or the distribution, has the same runtime as before, and only incurs a small cost of $\lambda\cdot\mathop{\mathrm{poly}}d$ in communication with the provers, where $\lambda$ is a bound on the allowed numerical precision. We also show how to handle arbitrary precision at the cost of incurring a tiny additive error, with essentially no overhead in communication and runtime.
@@ 1.1.3 Our techniques
Protocols for purely additive and for mixed errors.
As a warm-up, we describe our refereed learning protocols for the setting with additive error $\eta>0$ and zero-one metric on $\mathcal{Y}$. (We describe these protocols in more detail in Section 7.) The natural way for the learner to bound the additive error by $\eta$ without using the provers is to draw $O(1/\eta^{2})$ samples from $\mathcal{D}$}, query $f,h_{0},h_{1}$ at these points, and pick the hypothesis with the smaller empirical loss. When multiplicative loss $\alpha=1+\varepsilon$ with $\varepsilon>0$ is also allowed then the learner can choose to draw only $(1+\frac{1}{\varepsilon^{2}})\cdot\frac{1}{\eta}$ samples from $\mathcal{D}$}, find the set $S$ of samples on which $h_{0}$ and $h_{1}$ disagree, and then pick the hypothesis with the smaller empirical loss over the samples in $S$.
The learner can then use the provers as follows: instead of directly querying $f,h_{0},h_{1}$ on the sampled points, the learner can send the sample points to the provers and have the provers make the queries and report back the obtained values. If the provers disagree on any returned value, the learner picks one such value, makes the query itself, and proceeds with the values provided by the prover that reported the correct value.
@@
Next, consider the more challenging setting where only multiplicative error is allowed, i.e., $\eta=0$. Directly extending the above protocol from the mixed error case would be meaningless, since the number of samples required exceeds the domain size when $\eta$ approaches 0. An alternative approach would be to compute the empirical loss of the two hypotheses over a sufficiently large sample from the “disagreement set” $S={\left\{{x\mid h_{0}(x)\neq h_{1}(x)}\right\}}$; however, if the set $S$ is sparse then the learner cannot efficiently sample from it. Furthermore, it is not immediately clear how the learner can use the provers to provide it with correctly distributed samples from $S$. In particular, the above method of settling discrepancies between the provers regarding the value of either one of $h_{0},h_{1},f$ at a given point does not seem to be useful for the purpose of obtaining a random sample from $S$. To get around this issue we devise a certifiable uniform sampling protocol, described below, that allows the learner to obtain samples from $S$ that are guaranteed to be correctly distributed. Given this protocol, the learner simply picks the model with the smaller empirical loss over $O(1+\frac{1}{\varepsilon^{2}})$ random samples from $S$. Finally, the learner can offload all queries but one to the provers, as described earlier.
Certifiable uniform sampling.
In order to design a refereed learning protocol with a computationally efficient learner and bounded communication cost, we construct a protocol that allows the learner to efficiently generate, with the help of the provers, uniform samples from a set $S$ which can be both exponentially large and exponentially sparse. In particular, while the provers are assumed to know $S$, the learner does not need to compute $S$ itself.
As an aside, we note that the ability to use the two provers to sample from a distribution that the learner is unable to sample from by itself may be of more general interest beyond the current framework of refereed learning.
The certifiable sampling protocol uses two sub-protocols: (I) a protocol, Certifiable Sum, which allows the learner to efficiently obtain the size of $S$ with only a membership query oracle for $S$; (II) a protocol, Certifiable Index, which allows the learner to efficiently obtain the lexicographically ${i}^{\mathrm{th}}$ element in $S$. In each protocol, the learner runs in time $\mathop{\mathrm{poly}}d$. Now, to obtain $m$ uniformly random samples from $S$, the learner simply executes the first protocol to obtain $|S|$, then samples $m$ uniform indices $i_{1},\dots,i_{m}\in{\left\{{1,\ldots,|S|}\right\}}$, and executes the second protocol to obtain elements $S_{i_{1}},\dots,S_{i_{m}}$. These two sub-protocols are described next.
@@
This protocol allows the learner, assisted by provers, to evaluate quantities of the form $s=\sum_{x\in\{0,1\}^{d}}t(x)$, given only query access to the function $t$, in time that is polynomial in $d$. Now, to compute the size of the set $S$ we set $t$ to be the indicator function $t(x)=1$ if $x\in S$, and $t(x)=0$ otherwise.
@@
The certifiable index protocol.
This protocol allows the verifier to efficiently obtain the ${i}^{\mathrm{th}}$ element of a set $S\in\{0,1\}^{d}$ with ordering $\prec$, as long as the verifier has membership query access to $S$ and query access to the ordering $\prec$. At a high level, the verifier asks each prover for a candidate ${i}^{\mathrm{th}}$ element $x^{*}$ and then checks that $x^{*}$ is in $S$, and uses certifiable sum to confirm that there are i1i-1 elements in $S$ that are smaller than $x^{*}$. More formally, the verifier obtains the claimed ${i}^{\mathrm{th}}$ element $x^{*}$ from one of the provers, and then executes certifiable sum with the function $t(x)=\mathds{1}{\left[{x\in S\wedge x\prec x^{*}}\right]}$. If certifiable sum returns i1i-1 then the verifier outputs $x^{*}$ as the ${i}^{\mathrm{th}}$ element of $S$. Otherwise, it repeats the protocol with the other prover’s claimed ${i}^{\mathrm{th}}$ element. (Indeed, $x^{*}$ is the ${i}^{\mathrm{th}}$ element of $S$ if and only if there are exactly i1i-1 points $x$ such that $t(x)=1$.)
Extending to general sample distributions.
The protocol for certifiable uniform sampling is only relevant when the underlying distribution is the uniform distribution over $\{0,1\}^{d}$. We extend this protocol to handle sampling from arbitrary distributions $\mathcal{D}$} over $\{0,1\}^{d}$. This extended protocol allows the learner to efficiently generate certified samples from $\mathcal{D}$}, given query access to the distribution’s probability mass function $Q_{\mathcal{D}}$. This is done by using the certifiable sum protocol with a function $t$ that depends on the distribution via Birgé’s decomposition [Bir87]. (Birgé’s decomposition states that a a monotone distribution over $[N]$ can be approximated by a piecewise constant distribution over $O(\log N)$ many buckets.) See Lemma 3.1 for a formal statement of the certifiable sample protocol, and Theorem 4.2 for a formal statement of the resulting refereed learning protocol.

### v2

@@ Abstract
We initiate an investigation of learning tasks in a setting where the learner is given access to two competing provers, only one of which is honest. Specifically, we consider the power of such learners in assessing purported properties of opaque models. Following prior work in complexity theory that considers the power of competing provers in various settings, we call this setting refereed learning.
After formulating a general definition of refereed learning tasks, we show refereed learning protocols that obtain a level of accuracy that far exceeds what is obtainable at comparable cost without provers, or even with a single prover. We concentrate on the task of choosing the better one out of two black-box models, with respect to some ground truth. While we consider a range of parameters, perhaps our most notable result is in the high-precision range: For all $\varepsilon>0$ and ambient dimension $d$, our learner makes only one query to the ground truth function, communicates only $(1+\frac{1}{\varepsilon^{2}})\cdot\text{poly}(d)$ bits with the provers, and outputs a model whose loss is within a multiplicative factor of $(1+\varepsilon)$ of the best model’s loss. Obtaining comparable loss with a single prover would require the learner to access the ground truth at almost all of the points in the domain.
@@ 1 Introduction
Modern machine learning tasks require increasingly large amounts of data and computational power. As a result, model training has shifted from a task that almost anyone can perform on their own to a task that requires the assistance of external agents that are resource-abundant and have better access to the underlying data. Furthermore, one is often presented with models that purport to approximate some ground truth, but are not accompanied with a rigorous or trustworthy performance guarantee. Moreover, these models are often given as black-boxes—that is, properties of the model are difficult to assess given its description.
@@
This state of affairs naturally raises the need to verify claims of performance of such models, with significantly fewer resources than those needed to train comparable models and without fully trusting the parties making the claims. Verifying such claims appears hard: There is a rich literature on efficient verification of computations performed by powerful-but-untrusted parties (e.g., [Kilian92, Micali94, GoldwasserKR15]; see [WalfishB15] for a survey); however, such mechanisms appear ill-suited to the task of assessing a model’s accuracy, even if the code of the model is explicitly given. There is also a growing body of work on using powerful-but-untrusted intermediaries to learn properties of unknown ground-truth functions [GoldwasserRSY21, CanettiK21], as well as efficient verification of claims made by powerful provers regarding properties of huge combinatorial objects in general [ErgunKR04, RothblumVW13]. However, these works mostly focus on verifiable execution of a specific learning (or property testing) algorithm, rather than evaluating the performance of a given model without knowledge of the process used to train it. Perhaps closest to our task are the works of [HermanR22, HermanR24-1, HermanR24-2] that allow verifying claims about properties of distributions that are accessible only via obtaining samples. However, this is still a far cry from assessing performance of a model.
@@
Some natural properties of ML models can of course be assessed using standard methods. For instance, the loss of a given model w.r.t. some sample distribution and loss function can be approximated by computing its empirical loss on a large enough sample. However, this method can be prohibitively costly both in samples from the ground truth and in queries to the model. [GoldwasserRSY21] show a technique for pushing much of the burden to an external powerful-but-untrusted prover, but even this method incurs high cost: to obtain an additive bound of $\eta$ on the model’s error, the learner needs to both communicate $\eta^{-2}$ unlabeled samples to the prover and have the prover query and report the ground truth values at these points, and obtain $\eta^{-1}$ labeled samples which are hidden from the prover in order to verify the prover’s responses. To see that this is prohibitive, consider a setting where evaluating the ground truth requires performing an expensive physical experiment (e.g., as is required for validating the performance of AlphaFold [Jumper2021]); ideally, assessing the accuracy of a model should only require performing a small constant number of experiments.
@@
We would like to do better—in terms of the error, in terms of the access to the ground truth, and in terms of the communication with the prover. However, as evidenced by lower bounds proven in [GoldwasserRSY21], this appears hard—at least within the present framing of the problem.
@@ 1.1 This work
We show that the quality of black-box models can be assessed with significantly better accuracy and with significantly lower cost if the learner can interact with two powerful and competing provers. (The “competing” aspect of the provers can be captured as a zero-sum game where only one of the two provers wins; alternatively we can assume that one of the provers is honest.) The provers’ power can be manifested either in terms of their computational power, or in access to the ground truth, or in knowledge of the models, or any combination of these. This model is a natural extension of the refereed delegation of computation model of [FeigeST88, FeigeK97, CanettiRR11, CanettiRR13, KolR14] to our setting. Following the cue of these works, we coin the term refereed learning to denote the type of learning performed in this model. (See Section 1.2 for more discussion on and justification of this model.)
We first define refereed learning for a general setting. Next we focus on applying refereed learning to the following specific task, where we showcase the power of the refereed learning framework via concrete protocols. The learner–verifier [^1] We use the terms learner and verifier interchangeably. Indeed, the learner now doubles as a verifier. is presented with two candidate models that purport to compute the same ground truth, and is tasked with choosing the model that incurs the smaller overall loss with respect to some sample distribution and loss function. (The restriction to two candidates is not essential, but it helps make the model more concrete. It also facilitates envisioning each one of the provers as “trying to promote” a different one of the two models.)
The salient parameters we consider are (a) the overall loss of the output model relative to the better of the two competing models; (b) the number of learner queries to the ground truth and samples from it (both labeled and unlabeled ones); (c) the number of learner queries to the candidate models; (d) the computational complexity of the learner; (e) the computational complexity and query and sample complexity of the provers. We first sketch and briefly discuss definitions of refereed learning, then present our results, and finally discuss the new tools we develop to obtain these results.
A motivating example.
We focus on a setting where there is a known distribution over unlabeled samples, and where the ground truth is computable, but doing so is costly (e.g., it requires a physical experiment). The task is to compare the accuracy of two ML models (such as AlphaFold [Jumper2021]) where each model takes proteins as input and returns a prediction of how these proteins interact. The researcher (verifier) can specify a distribution over proteins of interest; however, assessing the quality of these predictions incurs high cost, even for a given distribution. Specifically, it often requires synthesis and purification of the actual proteins, followed by cryo-EM imaging to confirm the predicted protein interaction.
@@ 1.1.1 Defining refereed learning
We start with general refereed learning. Consider a learner-verifier $\mathcal{V}$ and two provers ${\mathcal{P}}_{0},{\mathcal{P}}_{1}$ that are presented with a ground truth function $f:\{0,1\}^{d}\rightarrow\mathcal{Y}$, a distribution $\mathcal{D}$} over $\{0,1\}^{d}$, and $k$ models (a.k.a. hypotheses) $h_{1},\ldots,h_{k}\in{\mathcal{H}}$, where $\mathcal{H}$} is some family of hypotheses. We will typically assume that $h_{1},\ldots,h_{k}$ and $f$ are accessed via queries and, depending on the setting, that $\mathcal{D}$} is accessed either via samples or via queries to its probability mass function, $Q_{\mathcal{D}}$, which maps $x\mapsto\operatorname*{{Pr}}_{X\sim{\mathcal{D}}}{\left[{X=x}\right]}$. To measure the learner’s performance we use a score function $\mathcal{S}$ (which assigns a score to each potential output of the learner, with respect to some sample distribution $\mathcal{D}$}, hypotheses $h_{1},\ldots,h_{k}$ and function $f$) and a target function $\mathcal{T}$}. In the definition below, $\alpha$ and $\eta$ respectively control the multiplicative and additive slack up to which the loss of the returned hypothesis matches that of the best hypothesis. We give protocols with purely multiplicative, purely additive, and mixed error guarantees.
@@
A protocol ${\mathcal{P}}_{0},{\mathcal{P}}_{1},\mathcal{V}]$ is an $(\alpha,\eta,\beta)$-refereed learning protocol, with respect to a family $\mathcal{H}$} of hypotheses, a score function $\mathcal{S}$ and target function $\mathcal{T}$}, if for all $b\in\{0,1\}$, $h_{1},\ldots,h_{k}\in{\mathcal{H}}$, and ${\mathcal{P}}^{*}_{1-b}$, the learner output $\rho$ satisfies
@@
For this general definition, we choose to minimize the score $\mathcal{S}$. This choice is motivated by the learning theory formulation of this problem, where the goal is to select the model (hypothesis) that minimizes a loss function quantifying the deviation of the hypothesis from some ground truth. Concretely, we focus on the setting where the learner selects between two models (hypotheses) $h_{0}$ and $h_{1}$, its output bit $\rho$ indicates the selected hypothesis, the general scoring function $\mathcal{S}$ measures the loss of the chosen model with respect to some metric $\ell$ on the domain $\mathcal{Y}$, and the target function $\mathcal{T}$} is the smaller of the losses of $h_{0}$ and $h_{1}$. (Additionally, we are primarily interested in the following loss: for some hypothesis $h$, sample distribution $\mathcal{D}$}, and ground truth $f$, let $\mathcal{L}_{{\mathcal{D}}}{\left({f,h\mid\ell}\right)}=\mathop{{\mathds{E}}\/}_{x\sim{\mathcal{D}}}{\left[{\ell{\big({f(x),h(x)}\big)}}\right]}$.) In this paper, we focus on refereed learning protocols satisfying the following definition.
@@
An alternative formulation would instead bound the expected score (respectively, loss). That is, the requirement would be that $\mathop{{\mathds{E}}\/}[\mathcal{S}(\rho,{\mathcal{D}},f,h_{1},\ldots,h_{k}))]\leq\alpha{\mathcal{T}}({\mathcal{D}},f,h_{1},\ldots,h_{k})+\eta$ (in the general case), or $\mathop{{\mathds{E}}\/}[\mathcal{L}_{\mathcal{D}}(f,h_{\rho}\mid\ell)]\leq\alpha\min_{s\in\{0,1\}}\mathcal{L}_{\mathcal{D}}(f,h_{s}\mid\ell)+\eta$ (in the case of minimizing the loss). While the two formulations are incomparable in general, our protocols satisfy both with similar parameters.
@@
The above definition posits that at least one of the provers follows the protocol. We note, however, that refereed learning protocols often seem to preserve their guarantees even when both provers are strategic with opposing goals. Some supporting evidence for the implication is the use of protocols developed in the refereed delegation of computation model in real-world applications where truth-telling is economically incentivized. Similar phenomena are manifested in the context of debate systems. See more details in Section 1.2.
@@ 1.1.2 Our results
We give protocols for the multiplicative, additive, and mixed error settings. In the additive error setting, we show how to use the two provers to obtain additive error similar to that obtained by the [GoldwasserRSY21] protocol mentioned above, while significantly reducing the learner’s interaction with both the ground truth and the models: this interaction now consists of only a single query.
@@
We focus primarily on designing protocols for the low-loss setting, which turns out to be significantly more challenging. Here we would like to guarantee that the learner makes the right choice even when the models’ losses are close to each other, up to a multiplicative factor of $1+\varepsilon$ for an arbitrarily small $\varepsilon>0$. Indeed, in such a setting, the number of samples needed to even observe the difference can be close to the entire sample space. Still, determining which of the two competing models incurs smaller loss can have significant ramifications in applications that require high precision (e.g., using ML models for medical predictions based on imaging or other multi-dimensional measurements, or for financial applications where even tiny error margins become significant over time). [^2] The purely multiplicative error setting can also be a better fit when working with competing provers. To see this, consider the following toy example. Suppose the verifier can only approximate provers’ additive errors up to some threshold $\eta$ (and is unable to distinguish provers’ losses that differ additively by at most $\eta$). Here each prover is only incentivized to learn up to loss $\eta$ if possible, but no better, since its hypothesis will be accepted with probability (at least) $\frac{1}{2}$ once it learns to loss at most $\eta$. Now suppose the verifier can approximate the ratio of the provers’ error up to $(1\pm\varepsilon)$. Here, a prover who can learn to loss 0 will do so (assuming sufficient incentives—e.g., no cost for additional computation).
@@
We then extend these results to handle arbitrary metric loss functions and arbitrary sample distributions. Our results also take into account the cost of obtaining the desired level of numerical precision. The protocol for this more general setting is guaranteed to select a model whose loss is at most a multiplicative factor of $3+\varepsilon$ worse than the best model. Moreover, the learner makes a single query to either $f$, the two models, or the distribution, has the same runtime as before, and only incurs a small cost of $\lambda\cdot\mathop{\mathrm{poly}}d$ in communication with the provers, where $\lambda$ roughly bounds the bit precision of the numerical representations used for sending values between parties in the protocol. We also show how to handle arbitrary precision at the cost of incurring a tiny additive error, with essentially no overhead in communication and runtime.
@@ 1.1.3 Our techniques
Protocols for purely additive and mixed errors.
As a warm-up, consider the setting with additive error $\eta>0$ and the zero-one metric on $\mathcal{Y}$. (We describe these protocols in more detail in Section 7.) The natural way for the learner to bound the additive error by $\eta$ without using the provers is to draw $O(1/\eta^{2})$ samples from $\mathcal{D}$}, query $f,h_{0},h_{1}$ at these points, and pick the hypothesis with the smaller empirical loss. When multiplicative loss $\alpha=1+\varepsilon$ with $\varepsilon>0$ is also allowed then the learner can choose to draw only $(1+\frac{1}{\varepsilon^{2}})\cdot\frac{1}{\eta}$ samples from $\mathcal{D}$}, find the set $S$ of samples on which $h_{0}$ and $h_{1}$ disagree, and then pick the hypothesis with the smaller empirical loss over the samples in $S$.
The learner can then use the provers as follows: instead of directly querying $f,h_{0},h_{1}$ on the sampled points, the learner can send the sampled points to the provers and have the provers make the queries and report back the obtained values. If the provers disagree on any returned value, the learner picks one such value, makes the query itself, and proceeds with the values provided by the prover that reported the correct value.
@@
Now consider the more challenging setting where only multiplicative error is allowed, i.e., $\eta=0$. Directly extending the above protocol from the mixed error case would be meaningless, since the number of samples required exceeds the domain size when $\eta$ approaches 0. An alternative approach would be to compute the empirical loss of the two hypotheses over a sufficiently large sample from the “disagreement set” $S={\left\{{x\mid h_{0}(x)\neq h_{1}(x)}\right\}}$; however, if the set $S$ is sparse then the learner cannot efficiently sample from it. Furthermore, it is not immediately clear how the learner can use the provers to provide it with correctly distributed samples from $S$. In particular, the above method of settling discrepancies between the provers regarding the value of either one of $h_{0},h_{1},f$ at a given point does not seem to be useful for the purpose of obtaining a random sample from $S$. To get around this issue we devise a certifiable sampling protocol, described below, that allows the learner to obtain samples from $S$ that are guaranteed to be correctly distributed. Given this protocol, the learner simply picks the model with the smaller empirical loss over $O(1+\frac{1}{\varepsilon^{2}})$ random samples from $S$. Finally, the learner can offload all queries but one to the provers, as described earlier. The certifiable sampling protocol makes heavy use of the certifiable sum protocol.
The certifiable sample protocol.
This protocol allows the verifier to efficiently obtain $m$ samples from an arbitrary distribution $\mathcal{D}$} over $\{0,1\}^{d}$—even when the support of $\mathcal{D}$} is both exponentially large and exponentially sparse. At a high level, the protocol does this through inverse CDF sampling facilitated by certifiable sum (described next). That is, the verifier specifies some total ordering on $\mathcal{D}$} and then samples a uniformly random value $p\in[0,1]$. The provers then return the element $x^{\prime}\in{\mathcal{D}}$ such that $\sum_{x\prec x^{\prime}}{\mathcal{D}}(x)<p\leq\sum_{x\preceq x^{\prime}}{\mathcal{D}}(x)$. The verifier then uses certifiable sum to confirm that these inequalities hold for the returned value $x^{\prime}$. See Lemma 3.1 for a formal statement of the certifiable sample protocol, and Theorem 4.2 for a formal statement of the resulting refereed learning protocol.
@@
This protocol allows the learner, assisted by provers, to evaluate quantities of the form $s=\sum_{x\in\{0,1\}^{d}}t(x)$, given only query access to the function $t$, in time that is polynomial in $d$.
@@
Theorem 1.3 (Refereeed learning for zero-one loss (Theorem 4.2, informal)).
There exists a protocol ${\mathcal{P}}_{0},{\mathcal{P}}_{1},\mathcal{V}]$ that, for all $\varepsilon,\beta>0$ and distributions $\mathcal{D}$} on $\{0,1\}^{d}$, is a $(1+\varepsilon,0,\beta)$-refereed learning protocol for $f$ and $\mathcal{D}$} with respect to $\ell_{\mathsf{zo}}$. The protocol has communication complexity and verifier runtime $\widetilde{O}{\left({(1+\frac{1}{\varepsilon})^{2}\cdot\mathop{\mathrm{poly}}d\cdot\log\frac{1}{\beta}}\right)}$, and the verifier makes 1 query to $f$.